OpenStudy (anonymous):

simplify the expression, log(10^6 x^2) a. 12log(x0 b.6x^2 c.12 d.6+2log(x)

OpenStudy (zarkon):

if it is \[\log_{10}(10^6x^2)\] then \[=\log_{10}(10^6)+\log_{10}(x^2)\] \[=6+2\cdot\log_{10}(x)\]